# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services

### v2

@@ VI-A Scaling and cost of the generation
TABLE VI: Average ($\pm$ 1-std) number of accumulated input & output tokens, generation time per test data generation for the SQL code generation services. One can estimate the actual monetary cost using a (external) pricing model for gemini-2.0-flash.
@@
Table VI presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model [35] (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
@@ VI-B Generating for test SQL query semantics
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query [36, 37]. While the NL2SQL prediction quality evaluation is beyond the scope of this study, a meaningful test data generation should provide a set of test data instances that are semantically relevant to the target SQL query. One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section IV-E.
1SELECT date_trunc(logdate, quarter), sum(conversions) / sum(clicks) 2FROM fake_table_1 3WHERE result_type = ’TEXT_AD’ AND logdate BETWEEN ’2022-01-01’ AND ’2022-12-31’ 4GROUP BY date_trunc(logdate, quarter);
Figure 11: Meaningful test data generation for this business analytics query requires generating data instances for each quarter in year 2022.
To illustrate the problem, consider a business analytic query in Figure 11. A meaningful test data for this query should ensure that enough example conversions instances are present across all the quarters in the requested date range. Our LLM-based generator can recognize such semantic importance of date_trunc(logdate, quater) and generate data for all quarters in the range. Moreover, if one has a domain knowledge about the data, such as “conversions were better in Q4 than Q1,” then it can be passed to the generator as user_criteria as part of input, which can guide the generation to favor Q4 instances. Such understanding and reasoning capabilities of LLMs can scale semantic-aware test data generation.
@@ VI-C Generation error analysis
  • Missing fields from large and complex schema, which requires some post-processing steps to identify and retry.
@@
1json[{ 2 "fake_service_id": 1234567890, 3 "user_dims": { 4 "dimension1": "value1", 5 "dimension2": 10 6 }, 7 "id": 12345, 8 "customer_dims": { 9 "dimension3": "value3", 10 "dimension4": 20 11 }, 12 "customer_details": { 13 "detail1": "detailValue1", 14 "detail2": 30 15 }, 16 "fake_service_dump_freshness_usec": 1678886400000000, 17 "user_details": { 18 "detail3": "detailValue3", 19 "detail4": 40 20 } 21 }, 22 ...]
@@
We have implemented some number of safeguards against LLM hallucination. Most notably, we have observed in some simple schema cases, the LLM introduced new fields – even nested ones. Figure 12 illustrates one such example where many more new fields are introduced next to just 3 required fields; the LLM even introduces nested fields.
The last item is very difficult to fully address. We use LLM-as-a-judge to validate the semantic coherence of generated data, which runs the risk of false positives, but we do not have an oracle (or access to the real production tables) to ensure all relevant values are covered and generated. If we can provide some domain knowledge, as discussed in Section VI-B, that could help inform the LLM to generate for particular cases. This is an important problem to further improve the fidelity of mock data generation in the production settings. There are also techniques to statistically impute missing values and estimate potential impact of missing values on query results [38, 39], which would be interesting to consider in the test data generation settings.
@@
In this paper, we addressed the critical challenge of generating high-fidelity and complex test data for Google SQL code generation services, particularly in environments where production data access is restricted. We highlighted the limitations of traditional data generation methods, focusing on their inability to handle large, deeply nested schemas and maintain semantic integrity to the target test cases. By leveraging the latest gemini-2.0-flash model, coupled with strategic pre- and post-processing, we demonstrated the capability to generate comprehensive test data instances compliant to large complex schemas that effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries. Our experiments demonstrate the practical effectiveness of our approach, delivering a reliable method for generating realistic test data essential in industrial environments where access to production datasets is restricted. While our implementation targeted Google SQL, the fundamental methodology is dialect-agnostic and should be readily applicable to other SQL code generation services. By deploying our test data generator, we’ve enabled more thorough SQL code testing, resulting in significantly improved test coverage.
As part of our future work, we plan to extend the capabilities of our framework to support other nested data types, such as XML, JSON, and Open API schemas. This will enable us to address a broader range of real-world applications beyond SQL code generation services, including RESTful API testing. Furthermore, we aim to explore techniques for enhancing the semantic coherence of generated data, ensuring that it not only adheres to structural constraints but also accurately reflects real-world business context and semantics. By expanding the scope and capabilities of our approach, we seek to provide a comprehensive solution for generating high-quality test data that addresses the evolving needs of industrial and large-scale data service development and testing.
Acknowledgment
We would like to thank John Cieslewicz, Kashmira Phalak, Yannis Papakonstantinou and the anonymous reviewers for their valuable feedback. We also thank Yuchen Zhang for his contribution in the early stages of this project.

### v3

@@ VI-A Scaling and cost of the generation
TABLE VII: Average ($\pm$ 1-std) number of accumulated input & output tokens, generation time per test data generation for the SQL code generation services. One can estimate the actual monetary cost using a (external) pricing model for gemini-2.0-flash.
@@
Table VII presents a cost analysis of the proposed test data generation measured using the ”gemini-2.0-flash” model. It quantifies the average and standard deviation of the accumulated number of input/output tokens, as well as the generation time per test query. The high variability in token counts and generation time suggests a wide range of query and schema complexities. We report the number of input and output tokens instead of the monetary cost directly, as it depends on the pricing model [36] (subject to change over time). The total cost of generation is proportional to the number of input/output tokens. Overall, our method can generate more comprehensive and complex test data (rows) much more efficiently than the manual generation – compared to the human labor cost, this is much cheaper and scalable, too.
@@ VI-B Generating for test SQL query semantics
Test data’s semantic coherence to its test target SQL query is critical in evaluating SQL code generation services, like NL2SQL. One of the key challenges in testing NL2SQL is evaluating semantic correctness of the predicted query [37, 38]. While the NL2SQL prediction quality evaluation is beyond the scope of this study, high-fidelity test data should not only be adhere to the underlying schema, but also semantically relevant to the target SQL query (natural language question). One way to address this is to cover all the edge cases for the relevant columns used in the SQL query, as we have focused in Section IV-E.
SELECT date_trunc(logdate, quarter), sum(conversions) / sum(clicks) FROM fake_table_1 WHERE result_type = ’TEXT_AD’ AND logdate BETWEEN ’2022-01-01’ AND ’2022-12-31’ GROUP BY date_trunc(logdate, quarter);
Figure 11: High-fidelity test data generation for this business analytics query requires generating data instances for each quarter in year 2022.
To illustrate the problem, consider a business analytic query in Figure 11. High-fidelity test data for this query should ensure that enough example conversions instances are present across all the quarters in the requested date range. Our LLM-based generator can recognize such semantic importance of date_trunc(logdate, quater) and generate data for all quarters in the range. Moreover, if one has a domain knowledge about the data, such as “conversions were better in Q4 than Q1,” then it can be passed to the generator as user_criteria as part of input, which can guide the generation to favor Q4 instances. Such understanding and reasoning capabilities of LLMs can scale semantic-aware test data generation.
@@ VI-C Generation error analysis
  • Missing fields from large and complex data format, which requires some post-processing steps to identify and retry.
@@
json[{ "fake_service_id": 1234567890, "user_dims": { "dimension1": "value1", "dimension2": 10 }, "id": 12345, "customer_dims": { "dimension3": "value3", "dimension4": 20 }, "customer_details": { "detail1": "detailValue1", "detail2": 30 }, "fake_service_dump_freshness_usec": 1678886400000000, "user_details": { "detail3": "detailValue3", "detail4": 40 } }, ...]
@@
We have implemented some number of safeguards against LLM hallucination. Most notably, we have observed in some cases with simpler data formats, the LLM introduced new fields – even nested ones. Figure 12 illustrates one such example where many more new fields are introduced next to just 3 required fields; the LLM even introduces nested fields.
Addressing the issue of missing edge cases relevant to specific SQL queries is challenging, as we lack an oracle (or access to production data). While our pipeline uses an ”LLM-as-a-judge” for semantic validation, this risks false positives. To improve fidelity, we could integrate domain knowledge (as discussed in Section VI-B) to guide the LLM toward specific cases. This remains an important problem for improving mock data generation in production. Furthermore, techniques for statistically imputing missing values and estimating their impact on query results [39, 40] warrant future investigation in this setting.
@@
In this paper, we addressed the critical challenge of generating semantically relevant and complex test data for Google SQL code generation services, particularly where production data access is restricted. We highlighted the limitations of traditional data generation methods that were used by developers, focusing on their inability to handle large, deeply nested data structures, like protocol buffers, and to maintain semantic integrity of the target test cases. By leveraging the latest Gemini 2.5 Flash model with strategic pre- and post-processing, we successfully generated comprehensive test data, compliant with large, complex data structures. This data effectively supports complex SQL testing, including queries with joins, aggregations, and nested subqueries, and significantly improves test coverage. Our experiments demonstrate a reliable and practical method for generating high-fidelity test data, which is essential for industrial environments where access to production data is restricted. While our current implementation targets Google SQL, the underlying methodology is dialect-agnostic.